Independant
How do you calculate probability given mean and standard deviation?
The mean and standard deviation do not, by themselves, provide enough information to calculate probability. You also need to know the distribution of the variable in question.
How do you calculate experimental probability?
Experimental probability is the number of times some particular outcome occurred divided by the number of trials conducted. For instance, if you threw a coin ten times and got heads seven times, you could say that the experimental probability of heads was 0.7.
Contrast this with theoretical probability, which is the (infinitely) long term probability that something will happen a certain way. The theoretical probability of throwing heads on a fair coin, for instance, is 0.5, but the experimental probability will only come close to that if you conduct a large number of trials.
What is the total area under the normal distribution curve?
The area under the normal distribution curve represents the probability of an event occurring that is normally distributed. So, the area under the entire normal distribution curve must be 1 (equal to 100%).
For example, if the mean (average) male height is 5'10" then there is a 50% chance that a randomly selected male will have a height that is below or exactly 5'10". This is because the area under the normal curve from the left hand side up to the mean consists of half of the entire area of the normal curve.
This leads us to the definitions of z-scores and standard deviations to represent how far along the normal curve a particular value is. We can calculate the likelihood of the value by finding the area under the normal curve to that point, usually by using a z-score cdf (cumulative density function) utility of a calculator or statistics software.
What is the probability of picking an ace of spades randomly from a pack of cards?
1/52 or one out of fifty-two
What is the key feature of probability sampling?
The key feature is that each sample of the given size has the same probability of being selected as the sample. Equivalently, each unit in the population has the same probability of being included in the sample.
What is the probability of rolling two odd numbers when two numbers are rolled?
3/6 chance of an odd number with one roll of the dice
and
3/6 for the second roll of a dice
so
3/6 x 3/6 = 9/36 = 3/12 = 1/4
there is a one in four chance that two odd numbers will be rolled from two roles of a dice.
What is the chance of picking a red king out of a pack of cards?
There are 52 cards in a pack. There are 2 Red kings (Heart and Diamond).
This means 2/52 or 1/26.
Why the sample should be normally distributed?
The sample should not be normally distributed.
If you have a population of size N from which a random sample of size n is to be drawn, then there are NCn possible samples. Each one of these must have the same probability of being thesample. That is, the sample is uniformly distributed - not Normally.
A fair coin is flipped 4 times what is the probability of getting more than 3 heads?
The probability of flipping a fair coin four times and getting four heads is 1 in 16, or 0.0625. That is simply the probability of one head (0.5) raised to the power of 4.
What are the number of outcomes when you roll two dice?
The number of possible outcomes when rolling two dice can be expressed through simple fractional multiplaction- There are 36 different options.
What is the probability of choosing an ace from a standard deck of playing cards?
4 in 52, or 1 in 13, or about 0.0769.
How many different arrangements of the letters in the word number can be made?
6! = 6x5x4x3x2x1 = 720 arrangements
1.16/z = -1.16/z
* * * * *
No!
Pr(Z > 1.16) = 0.123
So Pr(-1.16 < Z < 1.16) = 1 - 2*0.123 = 1 - 0.246 = 0.754
Because you are thinking permutations rather than combinations. There are four permutations of two coins, but there are only three combinations, because it does not matter which coin is heads and which coin is tails. As a result, the combination of heads and tails has a 0.5 probability, while two heads or two tails each have a 0.25 probability.
What is the role of an event official?
An Event official in Physical Education is a person who over sees everything.
How can you find the probability of two mutually exclusive events?
The calculation is equal to the sum of their probabilities less the probability of both events occuring.
If two events are mutually exclusive then the combined probability that one or the other will occur is simply the sum of their respective probabilities, because the chance of both occurring is by definition zero.
The probability is 5/9.